### v1

# CounselBench: A Large-Scale Expert Evaluation and Adversarial Benchmark of Large Language Models in Mental Health Counseling

Abstract
Large language models (LLMs) are increasingly proposed for use in mental health support, yet their behavior in realistic counseling scenarios remains largely untested. We introduce CounselBench, a large-scale benchmark developed with 100 mental health professionals to evaluate and stress-test LLMs in single-turn counseling. The first component, CounselBench-EVAL, contains 2,000 expert evaluations of responses from GPT-4, LLaMA 3, Gemini, and online human therapists to real patient questions. Each response is rated along six clinically grounded dimensions, with written rationales and span-level annotations. We find that LLMs often outperform online human therapists in perceived quality, but experts frequently flag their outputs for safety concerns such as unauthorized medical advice. Follow-up experiments show that LLM judges consistently overrate model responses and overlook safety issues identified by human experts. To probe failure modes more directly, we construct CounselBench-Adv, an adversarial dataset of 120 expert-authored counseling questions designed to trigger specific model issues. Evaluation across 2,880 responses from eight LLMs reveals consistent, model-specific failure patterns. Together, CounselBench establishes a clinically grounded framework for benchmarking and improving LLM behavior in high-stakes mental health settings.
[Uncaptioned image] Code & Data: https://github.com/llm-eval-mental-health/CounselBench
Disclaimer: this study does not endorse or recommend the use of LLMs in therapeutic practice. Our goal is to evaluate and characterize LLM behavior in counseling interactions for research only.
1 Introduction
Professional counseling is a critical pillar of mental healthcare, yet remains inaccessible to many due to structural barriers such as cost, limited provider availability, transportation challenges, and stigma DockeryEtAl2015 ; GargEtAl2022a ; Gary2005 ; TristianaEtAl2018 . Digital technologies have improved access through tools such as smartphone-based apps NearySchueller2018 , tele-therapy (telemedicine_mental_health, ) and online counseling services MarcelleEtAl2019 ; JoEtAl2023 , which allows users to seek support anonymously, affordably, and on demand NearySchueller2018 ; SagarPattanayak2015 ; ZhangEtAl2015 . However, even with these innovations, many individuals still lack adequate support due to difficualty finding a relatable therapist, distrust of the mental health systems NeukrugEtAl2017 ; JardineEtAl2024a , or a preference for peer-driven advice RepperandCarter2011 ; Gillard2019 . As a result, people often turn to general-purposed platforms like Reddit (mental_disorder_reddit, ), where the quality of advice is inconsistent and unregulated.
Recent advances in natural language processing (NLP) have introduced a new player in this space: large language models (LLMs). These models undergo large-scale pretraining on diverse data and have the ability to understand context and provide immediate answers, which presents a compelling opportunity to assist with mental health support chen2023soulchatimprovingllmsempathy ; Stade_Stirman_Ungar_Boland_Schwartz_Yaden_Sedoc_DeRubeis_Willer_Eichstaedt_2024_llm_change_behavioral_healthcare ; hua2024largelanguagemodelsmental . Individuals may engage directly with LLMs, such as ChatGPT, for psychological support (LLM_as_mh_resources, ) or through derivative applications specifically designed for mental healthcare (opportunities_risks_llm_mental_health, ).
Despite this rapid deployment, little is known about how LLMs behave in real-world counseling contexts. Recent incidents and critiques have raised concerns about inappropriate advice, lack of safeguards, and misalignment with clinical standards character_ai_lawsuit ; Duffy_2024 . Most existing medical NLP benchmarks focus on structured tasks like diagnosis or factual question answering galatzerlevy2023capabilitylargelanguagemodels ; nori2023capabilitiesgpt4medicalchallenge ; Xu_2024_mental_llm . However, effective counseling requires a different set of capabilities, such as emotional nuance, contextual sensitivity, and open-ended reasoning. Prior work on mental health dialogue has largely focused on simulated multi-turn conversations with small-scale human annotation wang2024clientcenteredassessmentllmtherapists ; iftikhar2024therapynlptaskpsychologists ; chaszczewicz2024multilevelfeedbackgenerationlarge ; wang2024patientpsiusinglargelanguage , leaving key gaps in evaluating how LLMs respond to authentic user concerns in a realistic setting.
In this work, we focus on a common and under-explored setting: single-turn counseling, where an individual asks a counseling question and receives a standalone response. This format reflects real-world use cases on online forums, counseling Q&A platforms, and chat-based interventions JAMA-patient-clinician-interaction2024 , and is particularly relevant for Just-in-Time Adaptive Interventions (JITAIs) just_in_time_adaptive_intervention . Because there is no opportunity for follow-up or repair, quality and safety in this setting are especially important.
We aim to answer three key questions:
  1. 1.
    How do LLM-generated counseling responses compare to those from online human therapists in perceived quality and safety?
  2. 2.
    Can LLMs reliably evaluate the quality of counseling responses?
  3. 3.
    What kinds of failure modes do LLMs exhibit in counseling settings, and can those failure modes be systematically elicited?
2 Overview of CounselBench
To address these questions, we construct CounselBench in collaboration with 100 mental health professionals. COUNSELBENCH is a two-part benchmark for evaluating LLMs in realistic, single-turn mental health counseling scenarios (Figure 1). It includes:
  • COUNSELBENCH-Eval: a benchmark of 2,000 expert evaluations of LLMs and online human therapist responses to real counseling questions, with written rationales and span-level annotations.
  • COUNSELBENCH-ADV: an adversarial dataset of 120 expert-authored questions designed to expose identified failure modes in LLMs.
Figure 1: Overview of CounselBench benchmark. COUNSELBENCH-EVAL (left) includes expert evaluation of LLMs and human responses to real counseling questions. COUNSELBENCH-ADV (right) includes adversarial questions authored by clinicians to target identified LLM failure modes. See Appendix B for distinct license/degree types and specialization areas.
Research Pipeline
We began by extracting 100 real counseling questions from a publicly available dataset scraped from CounselChat bertagnolli2020counselchat , covering 20 common mental health topics (§3.1). Each question was answered by GPT-4, Gemini, LLaMA-3, and an online human therapist (§3.2). To evaluate these responses, we defined six evidence-based dimensions grounded in counseling literature: overall quality, empathy, specificity, factual consistency, medical advice, and toxicity (§3.3).
We then recruited 100 mental health professionals, covering 43 categories of counseling expertise, to evaluate the quality and safety of each response (§3.4). Besides rating the messages using numeric scores, they were asked to provide span-level annotation and in-depth elaboration of their rating rationales, offering rich, contextual insights into how professionals formulate counseling messages deliberately. We then analyzed these expert evaluations to compare counseling quality across models and human responses (§3.5). The evaluations of human and model responses form the COUNSELBENCH-EVAL dataset. In parallel, we prompted eight LLMs to rate the same responses using the same criteria (§3.6).
To stress-test model safety further, we asked 10 experts to author 120 adversarial questions that are likely to surface the common LLM failure modes identified in COUNSELBENCH-EVAL. These questions were used to test eight LLMs under targeted issues and form the COUNSELBENCH-ADV dataset (§4). Together, CounselBench enables both broad benchmarking and fine-grained analysis of LLM performance, including comparison with human and automated evaluations.
3 CounselBench-EVAL: Benchmarking Counseling Quality with 100 Mental Health Professionals and 8 LLMs
We develop COUNSELBENCH-EVAL to systematically assess LLM behavior in realistic, single-turn counseling scenarios. This section describes the construction pipeline: question sourcing (§3.1), LLM response generation (§3.2), evaluation rubric design (§3.3), human annotation protocol (§3.4), results from expert scoring and analysis (§3.5), and experiments with LLM-based evaluators (§3.6).
3.1 Real-World Counseling Questions from Patients
Our study collects real-world counseling inquiries and responses between patients and clinicians from CounselChat, an online counseling platform where licensed therapists publicly respond to anonymous patient inquiries. We build on a version of the dataset compiled by bertagnolli2020counselchat , covering patient-clinician interactions posted before 2022. All online human responses were written by licensed human therapists before the widespread adoption of powerful LLMs like ChatGPT, which was released in late 2022.
To ensure topic diversity and response quality, we curated a diverse set of 100 counseling questions, with five questions in each of 20 topics: (1) depression, (2) relationships, (3) anxiety, (4) family conflict, (5) parenting, (6) self esteem, (7) relationship dissolution, (8) behavioral change, (9) anger management, (10) trauma, (11) marriage, (12) domestic violence, (13) grief and loss, (14) social relationships, (15) workplace relationships, (16) legal regulatory, (17) substance abuse, (18) counseling fundamentals, (19) eating disorders, (20) professional ethics. Full selection and filtering criteria are detailed in Appendix LABEL:appendix_qa_selection.
For each question, we selected the top upvoted therapist response as the "online human therapist" entry in our dataset. While upvotes are not a clinical quality measure, they offer a practical, publicly observable signal of relevance and perceived helpfulness in peer-supported mental health forums.
3.2 Generating Counseling Responses from LLMs
To ensure fair comparison and reduce annotator fatigue, we limited our evaluation to three LLMs per question in COUNSELBENCH-EVAL. Each mental health professional blindly annotated all four responses to the same question (one human-written and three LLM-generated), allowing consistent side-by-side assessment while keeping the task within a manageable cognitive load.
We selected three representative LLMs for evaluation: (1) GPT-4-0613 (via OpenAI API), a widely adopted commercial model, (2) LLaMA-3.3-70B-Instruct (accessed through Hugging Face), a leading open-source model, (3) Gemini-1.5-Pro (accessed via Google’s Generative AI SDK), which powers responses in Google Search and Assistant. Together, they represent major deployment paradigms: commercial APIs, open-source checkpoints, and search-integrated platforms. Each model is prompted with the original patient question and asked to generate a single-turn counseling response. Full prompt templates and generation details are provided in Appendix C.1.
3.3 Evaluation Rubric and Paradigms
To assess the quality and safety of counseling responses, we developed a multi-dimensional evaluation rubric based on clinical psychology literature and expert consultation. All metrics were rated using 5-point Likert scales ($1=$ the most negative; $5=$ the most positive) unless specified otherwise. Appendix A shows a screenshot of the rating questions and scales for evaluation.
Overall Quality captures the holistic judgment of each response.
Empathy measures whether the response demonstrates emotional attunement, compassion, or validation. Empathy is a central component of therapeutic modalities (e.g., person-centered therapy JaniEtAl2012 ), and is strongly linked to therapeutic alliance, a key predictor of positive therapy outcomes AckermanHilsenroth2003 .
Specificity evaluates how well the response is tailored to the user’s particular context, rather than offering generic or overly broad advice. Personalized replies help users feel heard and understood (ElkinandMcKay2014, ; StilesHorvath2017, ; KramerStiles2015, ), and contextual relevance is critical to effective support (HitchcockEtAl2024, ; StilesHorvath2017, ; KramerStiles2015, ).
Medical Advice flags whether the response includes therapeutic or diagnostic guidance that should only be provided by licensed professionals. Such advice typically requires individualized assessment and clinical expertise (FrankEtAl2020, ; LyonEtAl2011, ), and its timing and appropriateness must be carefully managed by trained practitioners (HillEtAl2023, ; PrassandKivlighanJr.2021, ). This dimension is treated as binary (Yes/No) with an additional “I am not sure" option, and higher rates of flagged content indicate safety concerns.
Factual Consistency checks whether the response aligns with accepted commonsense or clinical knowledge and avoids inaccurate or unsupported claims. Users often struggle to discern factual accuracy in online mental health content (Morahan-MartinAnderson2000, ). LLMs are known to hallucinate false information (SunEtAl2025, ), and even human counselors can produce factual errors in communication (Lobel2024, ). This dimension is rated on a 4-point scale, also with an additional “I’m not sure” option to accommodate cases where the rater cannot confidently assess it.
Toxicity assesses the presence of language that is potentially harmful, stigmatizing, dismissive, or ethically problematic. Individuals seeking mental health support are often in vulnerable states and may be particularly susceptible to damaging content (WebbEtAl2008, ; MitchellEtAl2017, ).
Rating Rationale & Text Extraction
To capture the context and reasoning behind human evaluations of counseling messages, we asked mental health professionals (§3.4) to provide open-text explanations for their ratings on Overall Quality, Medical Advice, Factual Consistency, and Toxicity. Additionally, we asked them to identify and extract specific portions of each counseling response that supported their ratings. Two researchers on the team checked and verified the quality of text responses after data collection. In general, human annotators provided detailed written feedback, with an average length of more than $600$ words per survey combining all rationales.
Figure 2: Distribution of (A) credential types and (B) counseling experience among the 100 annotators.
3.4 Expert Annotation Protocol
We implemented the evaluation rubric as a structured online questionnaire using Qualtrics, a wildly used platform for human subjects research. Prior to full deployment, we validated our questionnaire and study protocol with 8 participants through three rounds of pilot studies. Accordingly, we refined our evaluation questions and improved their clarity.
We recruited 100 licensed or professionally trained mental health practitioners across the U.S. through Upwork, the world’s largest talent platform commonly used in human-computer interaction (HCI) research for sourcing domain-expert annotators cscw-gig-economy-Upwork . All annotators met the requirement of having formal clinical counseling training or certification. Collectively, the annotators held 32 distinct types of licenses and degrees, and reported expertise spanning 43 specialized counseling areas; see Appendix B for more details. This broad coverage ensures that our evaluation reflects a diverse and representative cross-section of real-world counseling practice. Their credentials and clinical experience are summarized in Figure 2; additional demographics information is provided in Appendix B. Our annotator sample reflects the current demographics of the counseling profession in the U.S., which is predominantly composed of White and female practitioners McGuireMiranda2008 ; Verdieu2024 ; APA22 .
Each annotator was assigned a personalized survey containing 5 counseling questions, each paired with 4 responses: one written by an online human therapist and three generated by different LLMs. This resulted in 20 QA pairs per annotator. Each pair was independently evaluated by five different professionals, supporting inter-rater reliability and reducing individual bias. Importantly, for each question, all four responses were rated by the same group of annotators, enabling direct and fair comparison across models. In total, the study includes 2,000 annotated QA pairs.
Annotators were blind to the source of each response and were not informed that some had been generated by LLMs. Surveys were randomly sampled and individually assigned. Each annotator spent around 40 minutes per survey and was compensated \$20-\$35 depending on professional credentials and experience. This study was reviewed and approved as exempt research by the Institutional Review Board of University of Southern California (Study ID: UP-25-00022).
3.5 Expert Evaluation Results: Counseling Quality Across LLMs and Humans
We present the results of expert annotations comparing the counseling responses generated by GPT-4, LLaMA-3.3, Gemini-1.5-Pro, and online human therapists.
Figure 3: Average expert ratings.
Overall performance
Table 1 reports the average expert-annotated scores for each model, and Figure 3 provides a visual summary. Across most major evaluation dimensions, LLM-generated responses significantly outperformed those written by online human therapists. Among the models, LLaMA-3.3 received the highest overall ratings, leading on five of six dimensions. To assess significance, we performed Wilcoxon signed-rank tests wilcoxon_test_2005 , comparing the expert ratings of each LLM’s responses to those of online human therapists. We verified the consistency of our expert annotators’ ratings by calculating inter-rater reliability; annotator responses across measured metrics yielded Krippendorff’s alpha values greater than $0.7$, indicating a high level of inter-rater agreement inter-rater-reliability_2016 (See Table 2).
Table 1: Average expert ratings of counseling responses across six evaluation criteria. Each response was rated by five mental health professionals; scores were first averaged by question, then by model. Responses marked "I am not sure" were excluded. For Medical Advice, the percentage of “Yes” responses was computed per question (excluding “I’m not sure”) and averaged over all questions. Statistical significance was assessed using Wilcoxon signed-rank tests wilcoxon_test_2005 , comparing scores (averaged per question) between each LLM and online human therapists. Significance thresholds: $p^{*}<0.05$, $p^{**}<0.01$, $p^{***}<0.001$. Significance was not computed for Medical Advice, which was treated as a categorical variable.
Overall $\uparrow$ (range:1-5) Empathy $\uparrow$ (range: 1-5) Specificity $\uparrow$ (range: 1-5) Medical Advice (percentage of “Yes") Factual Consistency $\uparrow$ (range: 1-4) Toxicity $\downarrow$ (range: 1-5) GPT4     $3.28^{***}$     $3.37^{***}$     $3.46^{n.s.}$      0.07      $3.53^{***}$     $1.78^{***}$ Llama 3.3     $4.29^{***}$     $4.22^{***}$     $4.63^{***}$      0.14      $3.70^{***}$     $1.36^{***}$ Gemini-1.5-Pro     $3.26^{***}$     $2.76^{n.s.}$     $3.50^{n.s.}$      0.08      $3.52^{***}$     $1.64^{***}$ Online Human Therapists     2.60     2.72     3.29      0.17      2.92     2.56
Table 2: Averaged (over questions) Krippendorff’s alpha with ordinal weighting. Krippendorff’s alpha is not applied to binary Medical Advice.
Overall Score Empathy Specificity Factual Consistency Toxicity Krippendorff’s alpha (K-$\alpha$) 0.82 0.83 0.82 0.75 0.72
Despite strong scores, a higher proportion of LLaMA-3.3’s outputs (14%) were flagged for providing unauthorized medical advice (e.g., recommending therapy techniques), highlighting an important safety concern. In contrast, GPT-4 responses were more likely to include safety disclaimers, with approximately one-third of outputs explicitly declining to answer and recommending consultation with human professionals.
Qualitative Analysis of Annotator Rationales
To better understand expert judgments, we conducted a thematic analysis following established protocol Clarke_Braun_thematic_analysis_2017 . Two researchers on our team independently reviewed and thematically coded annotators’ open-text explanations. Seven primary issues among low-quality counseling responses are identified: 1) Lacking empathy or emotional attunement, 2) Displaying an inappropriate tone or attitude (e.g., dismissive, superficial), 3) Providing inaccurate suggestions (e.g., containing wrong information or making recommendations without sufficient evidence), 4) Offering unconstructive feedback (e.g., lacking clarity or actionability), 5) Demonstrating little personalization or relevance, 6) Containing language or terminology issues (e.g., typos, grammatical errors), and 7) Overgeneralizing or making judgments and assumptions without sufficient context.
We then prompted GPT-4.1 to label all low-rated responses (overall score $\leq$ 2) using these categories. Each response could exhibit multiple issues. A detailed breakdown is provided in Figure LABEL:fig:low_overall (Appendix LABEL:sub_appendeix_llm_judge_eval). The top two most frequently flagged failure types for each model are:
  • GPT-4: Offering unconstructive feedback (e.g., lacking clarity or actionability) (49.3%) and demonstrating little personalization or relevance (41.3%).
  • Llama-3.3: Overgeneralizing or making judgments and assumptions without sufficient context (66.7%) and offering unconstructive feedback (e.g., lacking clarity or actionability) (27.8%).
  • Gemini-1.5-Pro: Lacking empathy or emotional attunement (44.1%) and offering unconstructive feedback (e.g., lacking clarity or actionability) (40.7%).
  • Online human therapist: Overgeneralizing or making judgments and assumptions without sufficient context (46.7%) and displaying an inappropriate tone or attitude (e.g., dismissive, superficial) (36.1%).
Clinical Boundaries: Unauthorized Medical Advice in LLM responses
To further examine safety risks, we analyzed the response excerpts flagged by expert annotators for containing medical advice. The most concerning identified issue was the inclusion of unauthorized clinical recommendations, especially prescribed psychotropic treatments (i.e., medications for mental health treatments, according to National Institute of Mental Health mental-health-medication-NIMH ). Among the LLMs, GPT-4 occasionally recommended specific medications, such as selective serotonin reuptake inhibitors (SSRIs). Annotators also identified that all three LLMs occasionally advised therapy techniques like cognitive behavioral therapy (CBT) or mindfulness. While these practices are common in clinical care, professionals emphasized that they require tailored application and oversight.
3.6 Can LLMs Reliably Judge Counseling Quality?
Human expert evaluation provides gold-standard insight but is costly and difficult to scale gehrmann-etal-2021-gem . As LLMs are increasingly considered for deployment in mental health applications, one important question is whether models can serve as their own evaluators. Prior work has proposed the “LLM-as-Judge” paradigm for tasks such as summarization and factuality madaan2023selfrefine ; manakul2023selfcheckgpt , but its reliability in high-stakes, subjective domains like counseling remains unclear.
To investigate this, we tested eight advanced LLMs as automated judges, as listed in Figure 4. Each model was prompted with the same QA pairs used in Section 3.4, along with the evaluation criteria provided to human experts. Prompts followed the survey format used in our human annotation protocol, with minor formatting constraints tailored to each model; see Appendix LABEL:sub_appendeix_llm_judge_eval for details.
Figure 4: Average evaluation scores across six dimensions (subplots) for counseling responses generated by GPT-4, LLaMA-3.3, Gemini-1.5-Pro, and online human therapists (x-axis in each subplot). Each colored line represents one evaluator, including eight LLM-based judges and human experts (red). Higher values indicate better performance except for Toxicity and Medical Advice. See Table LABEL:llm_judge_numeric_score for full numerical results.
LLM Judges Overrate Responses and Miss Safety Failures. Figure 4 presents the scores provided by LLM judges, compared with human expert annotations. LLM-based evaluations consistently yielded inflated scores relative to human ratings. This was most pronounced for Factual Consistency, where LLM judges frequently assigned perfect or near-perfect scores to LLM-generated responses, regardless of their actual content. Ratings for Toxicity showed minimal variance across all sources. LLM judges uniformly assigned the lowest toxicity scores, even when human experts had flagged content as potentially harmful or inappropriate. This lack of sensitivity indicates a critical weakness in the use of LLMs for safety evaluations in high-risk domains.
LLM Rankings Diverge from Human Preferences.
Beyond absolute scores, we compared model rankings based on average quality, as shown in Figure 4 and Table LABEL:llm_judge_numeric_score. The preferences expressed by LLM judges diverged sharply from those of human annotators. Most notably, Gemini-1.5-Pro was rated as the lowest-performing model by human expert annotators but was consistently ranked above GPT-4 by every LLM judge in terms of Overall Quality. This calls into question the reliability of LLMs in identifying relative quality, even under matched evaluation instructions.
LLM Judges Rarely Flag Problematic Text.
We also tested whether LLM judges could identify problematic content at the sentence level. As in the human evaluation protocol, each model was asked to extract spans containing medical advice, factual errors, or toxic language. We compared the extracted text against sentences flagged by at least two human expert annotators, treating these as widely agreed indicators of failure.
To standardize comparison, we mapped all annotations to complete sentences; see Appendix LABEL:sec:sentence_level for details. Table 3 reports the proportion of widely-flagged sentences detected by each LLM judge. Most LLMs failed to detect even a single flagged toxic or incorrect sentence. Some models detected limited instances of unauthorized medical advice, but none showed robust or reliable detection across safety categories. These results help explain the inflated consistency and toxicity scores in earlier evaluations and highlight the limitations of LLM-based judging in high-stakes settings.
Table 3: Percentage of widely flagged sentences (labeled by at least two annotators) detected by each LLM judge. Rows correspond to LLM judges, and columns correspond to four response sources: GPT-4, LLaMA-3.3, Gemini-1.5-Pro, and online human therapists. Higher values reflect closer alignment with expert annotations.
Flagged Medical Advice Flagged Factual Errors Flagged Toxic Content Judges GPT-4 Llama-3 Gemini Human GPT-4 Llama-3 Gemini Human GPT-4 Llama-3 Gemini Human GPT-3.5-turbo 0 0 0 0 0 0 9.1% 8% 0 0 10% 2.4% GPT-4 25% 11.1% 11.1% 23.1% 0 0 0 5.3% 0 0 0 0 Llama-3.1 37.5% 5.6% 22.2% 30.8% 0 0 0 8% 0 0 0 0 Llama-3.3 37.5% 11.1% 33.3% 34.6% 0 0 9.1% 10.7% 0 0 0 0 Claude-3.5 25% 11.1% 11.1% 3.9% 0 0 0 6.7% 0 0 0 0 Claude-3.7 37.5% 11.1% 33.3% 7.7% 0 0 9.1% 8% 0 0 0 0 Gemini-1.5-Pro 12.5% 5.6% 33.3% 15.4% 0 0 0 61.3% 0 0 0 0 Gemini-2.0-Flash 37.5% 11.1% 11.1% 26.9% 0 0 0 6.7% 0 0 0 0
4 CounselBench-ADV: An Adversarial Benchmark for Surfacing LLM Failures in Counseling
While expert annotators rated LLM-generated counseling responses as generally high quality, our in-depth analysis revealed recurring issues, such as inappropriate advice or generic feedback (§3.5). To probe these failure modes more systematically, we introduce CounselBench-ADV, an adversarial benchmark constructed by mental health professionals.
Identifying Fine-grained Failure Modes
To construct a targeted adversarial benchmark, we first need to isolate concrete failure modes that LLMs exhibit in realistic counseling settings. Broad categories (e.g., "lacking empathy" and "lacking personalization") identified in Section 3.5 are too general to support targeted adversarial construction. To enable precise probing, we conducted a deeper review of expert rationales and flagged responses to extract six concrete failure modes:
  • GPT-4: The response provides specific medication (1. medication) and suggests specific therapy techniques suggestions (2. therapy).
  • Llama-3.3: The response speculates about medical symptoms (3. symptoms) and is judgmental (4. judmental).
  • Gemini-1.5-Pro: The response is apathetic (5. apathetic) and is based on unsupported assumptions (6. assumptions).
Here we selected two per model that best represent the kinds of specific issues each model; see Table LABEL:tab:six_issues for details. Each issue is linked to real examples from CounselBench-EVAL, providing clear foundations for clinicians to construct new questions that elicit these failures in a controlled and measurable way.
Adversarial Dataset Construction
We rehired 10 mental health professionals who previously participated as expert annotators in CounselBench-EVAL. Each clinician was provided with detailed definitions and real examples illustrating each failure type, drawn directly from flagged responses in CounselBench-EVAL (Table LABEL:tab:six_issues). They were instructed to independently author two realistic, single-turn counseling questions for each failure mode, i.e., crafted in such a way that it would be plausible for an LLM to fall into the corresponding error.
Each clinician authored 12 adversarial questions (2 per issue × 6 issues), resulting in a total of 120 adversarial questions. Clinicians were asked to ensure the realism and diversity of the questions, reflecting what actual patient might ask during counseling. Importantly, these questions do not directly contain the failure, but are designed to trigger the failure in LLM responses. The resulting dataset provides a high-precision probe of specific vulnerabilities in LLM-generated counseling.
Eliciting and Evaluating Model Failures
To evaluate model behavior under adversarial conditions, we used the same prompting setup as in CounselBench-EVAL3.2) to generate counseling responses for each of the 120 adversarial questions. Each prompt was submitted to eight LLMs in Table 4. For each model, we issued the same prompt three times to sample variations in generation. In total, this yielded 2,880 QA pairs (120 questions × 8 models × 3 generations).
We used a hybrid annotation protocol to determine whether each adversarial question successfully triggered its target failure mode. One mental health professional reviewed a random subset of 96 QA pairs (3% of the full set), while GPT-4.1 was used to annotate the full set. We provided GPT-4.1 with detailed definitions of each failure type and included an in-context example drawn from CounselBench-EVALLABEL:sec:aversarial_by_gpt_41). In this way, the GPT evaluator had both conceptual clarity and concrete guidance to support consistent and accurate labeling. Agreement between the expert and GPT annotator reached 72.9%, indicating high reliability inter-rater-reliability_2016 .
Table 4: Fraction of responses identified to contain each targeted failure mode. Higher values reflect greater model vulnerability to the targeted issue.
Specific Issue GPT-3.5-Turbo GPT-4 Llama-3.1 Llama-3.3 Claude-3.5-Sonnet Claude-3.7-Sonnet Gemini-1.5-Pro Gemini-2.0-Flash 1. Medication   0.25   0.32   0.40   0.55   0.47   0.53   0.42   0.47 2. Therapy   0.28   0.25   0.60   0.53   0.50   0.53   0.30   0.35 3. Symptoms   0.20   0.82   0.83   0.83   0.87   0.83   0.67   0.68 4. Judgmental   0.12   0.02   0.22   0.22   0.15   0.08   0.15   0.08 5. Apathetic   0.85   0.22   0.18   0.13   0.10   0.12   0.47   0.42 6. Assumptions   0.53   0.35   0.73   0.80   0.57   0.32   0.42   0.48
Adversarial Questions Effectively Trigger Targeted Failures
Table 4 reports the frequency of each failure mode across model outputs based on GPT-4.1 evaluation. The adversarial questions effectively elicited the intended issues in a large fraction of cases. Notably, speculation about medical symptoms emerged as the most prevalent failure: it was triggered in 82% of GPT-4 responses, 83% of LLaMA series responses, over 83% of Claude models, and more than 67% of Gemini outputs. This pattern raises serious concerns, as such speculation can lead to inappropriate clinical guidance without proper diagnosis. In contrast, judgmental tone was the least common issues across models, appearing in only 2% of GPT-4 responses and 22% of responses from the LLaMA series.
Interestingly, the failure modes revealed through CounselBench-ADV are not randomly distributed across models, but exhibit clear family-level patterns. Models within the same family (e.g. LLaMA, Gemini, and Claude) tend to exibit similar distributions across the six failure modes, while GPT-3.5-Turbo and GPT-4 show distinct patterns. This suggests that failure modes remain relatively consistent within model families, but can shift substantially across major model upgrades (openai2024gpt4technicalreport, ).
5 Related Work
A growing body of work has attempted to construct datasets for evaluating counseling quality. Foundational efforts by Pérez-Rosas et al. perez-rosas-etal-ACL2019-good-counselor introduced one of the first annotated datasets for motivational interviews (MI), distinguishing high- and low-quality sessions based on linguistic and behavioral features. Many of these early datasets were built by extracting and transcribing dialogues from actual counseling sessions between patients and counselors (AnnoMI-ICASSP-2022, ; CUEMPATHY-ISCSLP2022, ; SPARTA-HOPE-data, ; PAIR-Dataset-EMNLP-2022, ). Other datasets focused on detecting markers of mental health disorders such as depression or anxiety MOTIVATE-IJCNN ; gratch-etal-2014-DAIC-WoZ-Dataset ; mentalchat16k , annotating student counseling exchanges ConvCounsel-student-counseling-2024 , or capturing specific therapeutic constructs such as empathy MEDIC-MM2023 ; CUEMPATHY-ISCSLP2022 . To mitigate privacy concerns associated with clinical data, some studies turned to social media ALONE-Dataset or summarized clinical notes MEMO-KDD2022 ; others used LLMs to simulate counseling content ict4awe25 ; BOLT-Chiu-Althoff-2024 . Across these works, scholars recognized specific challenges on dataset construction:
Grounding Evaluation in Clinical Expertise
Most prior work adopts evaluation criteria from psychology literature li-etal-2024-understanding-therapeutic ; chiu2024computationalframeworkbehavioralassessment , but rarely involves clinicians in rubric design. As a result, many rely on untrained crowdworkers or subjective heuristics Young_AI_human_response_2024_chi . In contrast, we develop our evaluation protocol collaboratively with clinical psychologists and scale annotation to 100 licensed or trained professionals, ensuring clinical validity and reproducibility.
Counseling Format: Single-turn vs. Multi-turn.
Many existing datasets rely on multi-turn dialogues from real counseling sessions (wang2024clientcenteredassessmentllmtherapists, ; iftikhar2024therapynlptaskpsychologists, ), which, while realistic, introduce challenges for evaluation. For instance, nonverbal cues are lost in transcripts AnnoMI-ICASSP-2022 ; ConvCounsel-student-counseling-2024 ; SPARTA-HOPE-data , and entangling memory, turn-taking, and contextual reasoning can obscure the source of model failures. Moreover, session-specific nuances often hinder generalization. We instead focus on the under-explored single-turn setting, prevalent on public counseling platforms (e.g., CounselChat), where each user question receives a standalone answer. This design facilitates fine-grained, interpretable evaluations and allows direct comparison between models and humans.
LLM-Assisted Annotation and Simulation.
Recent work has leveraged LLMs both to simulate therapy content BOLT-Chiu-Althoff-2024 ; ict4awe25 and to construct artificial patients Scaffolding_Empathy_patient_sim_2025 or therapists Wang_sim_therapists_2024 . LLMs have also been applied as automated evaluators of counseling responses li-etal-2024-understanding-therapeutic ; yosef-etal-2024-assessing ; zhao2024escevalevaluatingemotionsupport , with some studies showing that modern models encode useful therapeutic knowledge ConceptPsy_2025 . However, most of these efforts relied on very limited expert validation. Our work addresses this gap by systematically comparing LLM-generated evaluations against scores from 100 mental health professionals across 2,000 responses. This constitutes the most comprehensive comparison to date between LLM and human judgments in mental health counseling.
6 Limitations, Conclusion, & Future Work
This work introduces a six-dimensional evaluation paradigm for single-turn counseling responses and publishes a corpus of 2,000 expert-annotated evaluations, coupling numeric ratings with in-depth written rationales (CounselBench-EVAL), and 120 clinician-designed adversarial prompts (CounselBench-ADV). To our knowledge, CounselBench is the largest expert-annotated benchmark in this domain. It is the first to encompass both the expertise and the scale of annotators. Our results show that state-of-the-art LLMs outperform community up-voted responses written by online therapists on all quality dimensions, yet our analysis also uncovers persistent pitfalls when LLMs were used either to generate or evaluate counseling responses.
We acknowledge that the current work entails several limitations despite our best efforts. First, although the human-written responses were top-rated on online counseling platforms, they were produced with limited patient information and no in-person contact, which might not represent the best quality of clinician-provided counseling. In fact, clinical psychology literature has long recognized that even well-crafted counseling responses may carry structural biases Smith_counseling_bias_1980 ; clinical-law-review-biases-2002 ; Morrow_Deidan_counseling_bias_1992 ; this highlights the need to diversify comparative standards for evaluating LLM-generated counseling quality. Relatedly, although we aimed to align our expert annotator demographics with real-world trends in the counseling profession, the field itself has long exhibited demographic imbalances.
Alongside addressing these limitations, overcoming current model shortages will require aligning evaluation models more closely with expert standards, for example, through fine-tuning or reinforcement learning with human feedback (RLHF). Expanding the taxonomy of models’ failure modes and integrating them into training and assessment is another pressing direction. Future work could also extend the evaluation paradigm to multi-turn dialogues to see how risks evolve as context accumulates, and replicate the study across languages, clinical domains, and geolocations to test the generality of our current case studies while examining the utility of CounselBench for developing more robust models for clinical use. Together, these efforts will move the field toward LLMs that are not only capable but consistently safe and trustworthy in mental health applications.
Acknowledgments
We thank the mental health professionals who contributed to this project. Their careful annotations, clinical insights, and generous support were invaluable to our study. We also thank Athan Li, Konghao Zhao, and Shupeng Cheng for participating in the initial survey rounds and offering helpful input, and Maya Andres for reviewing and providing feedback on our early survey materials. Finally, we thank Bingsheng Yao and Dakuo Wang for their helpful discussions and feedback on COUNSELBENCH-EVAL.
Appendix A Survey Design
We utilized Qualtrics to conduct our surveys. We built a template as shown in Figure 5 and then populated surveys for all question and response pairs. Each annotator evaluated five randomly sampled questions, each accompanied by four answers — one from GPT-4, Llama-3.3, Gemini-1.5-Pro, and an online human therapist. For each post-response pair, annotators evaluated across our six metrics, providing rationals and span-level annotations.
Figure 5: Survey interface: annotators read a user post and one response (left) and rate the response on six criteria with Likert scales and text-evidence boxes (right).
Appendix B Annotator Demographic Overview
This section summarizes annotator demographics. Table LABEL:tab:license_degree_summary reports the distribution of counseling licenses and degree types, while Table 5 lists the specialization areas of the 100 mental-health professionals. Figure 6 shows the ethnicity and gender distribution of annotators.
Figure 6: Ethnicity (left) and gender distribution (right) of the 100 mental-health professional annotators.
Table 5: Distribution by Counseling Specialization. Annotators may have multiple specializations and thus may be counted more than once in this table.
Counseling Specialization Count
Depression/anxiety    77
Trauma/PTSD    62
Child/adolescent therapy    40
Crisis intervention    33
Grief counseling    23
Substance use disorders    22
Couples/family therapy    21
Obsessive-Compulsive Disorder (OCD)    4
Attention-Deficit/Hyperactivity Disorder (ADHD)    3
LGBTQ + LGBTQIA+ issues    3
Prefer not to answer    2
Autism Spectrum Disorder (ASD)    2
Relationships    2
Life transitions    2
Self-Esteem    2
School counseling    2
Perinatal mental health    2
Sexual Disorders    1
Dissociation    1
Bipolar    1
Holistic wellness    1
Chronic illness/pain    1
Use substances issues    1
Narcissistic abuse    1
Codependency    1
Gender    1
Sexual orientation    1
Young adults Eating disorders    1
Women’s issues    1
Geropsychology    1
Neurodiversity    1
Exposure and Response Prevention(ERP)    1
Cognitive Behavioral Therapy(CBT)    1
Eating disorder/body dysmophia/HAES counseling    1
Religious trauma    1
Relational concerns    1
Identity development    1
Neuro-affirming mental health therapy    1
Disabilities    1
Corrections    1
Disordered eating    1
Other diverse identities presented in multicultural counseling    1
Psychosis    1
Mood disorders    1
Table 6: Annotator Degrees, Licenses, and Additional Certifications. Annotators may hold multiple credentials and may therefore appear more than once in the table.
Licensed Professionals Count
Licensed Professional Counselor (LPC)     25
Licensed Clinical Social Worker (LCSW)     18
Licensed Marriage and Family Therapist (LMFT)     8
Licensed Mental Health Counselor (LMHC)     5
Licensed Addiction Counselor (LAC)     2
Licensed Independent Social Worker of Clinical Practice (LISW-CP)     1
Licensed Clinical Professional Counselor (LCPC)     1
Clinical Supervisor (CS)     1
Licensed Addiction Counselor Supervisor (LACS)     1
Doctoral-Level Degrees
Psychologist (Ph.D./Psy.D.)     10
Psychiatrist (MD)     1
Trainees / Associates / Under Supervision
Licensed Associate Professional Counselor (LAPC)     3
Licensed Master Social Worker (LMSW)     2
Licensed Professional Counselor Associate (LPCA)     1
Licensed Mental Health Counselor Associate (LMHCA)     1
Licensed Associate Counselor (LAC)     1
Licensed Professional Counselor Candidate (LPCC)     1
Licensed Clinical Social Worker Associate (LCSWA)     1
Licensed Psychological Associate (LPA)     1
Licensed Social Worker (LSW)     1
Licensed Alcohol and Drug Counselor (LADC)     1
Currently in training/supervision    11
Other Degrees
Education Specialist Degree in school psychology     2
Master’s Degree in Counseling     1
Master’s Degree in School Counseling     1
Master’s Degree in Counseling Psychology     1
Master’s Degree in Psychology     1
B.A. in psych    1
Formerly Licensed / Associate
Licensed Associate Counselor (LAC)    1
Licensed Marriage and Family Therapist (LMFT)    1
Counseling-Related Certifications
National Certified Counselor (NCC)     2
School Psychologist License    1
K-12 School Counselor certification    1
Certified in FFT (Functional Family Therapy)    1
Trauma Informed CBT Certified    1
Appendix C Experiment Details
C.1 Response Generation
We use a zero-shot prompting setup with standardized instructions across models. Decoding parameters are fixed: temperature = 0.7, top-p = 1.0, and a maximum output length of 1024 tokens. To ensure consistency with human responses, we apply rejection sampling to enforce a 250-word limit, regenerating outputs as needed.
We use the following prompt to the LLM to generate a response for each user question, and then provide the original user query directly to generate a response.