### v3

# CounselBench: A Large-Scale Expert Evaluation and Adversarial Benchmarking of Large Language Models in Mental Health Question Answering

Abstract
Medical question answering (QA) benchmarks often focus on multiple-choice or fact-based tasks, leaving open-ended answers to real patient questions underexplored. This gap is particularly critical in mental health, where patient questions often mix symptoms, treatment concerns, and emotional needs, requiring answers that balance clinical caution with contextual sensitivity. We present CounselBench, a large-scale benchmark developed with 100 mental health professionals to evaluate and stress-test large language models (LLMs) in realistic help-seeking scenarios. The first component, CounselBench-EVAL, contains 2,000 expert evaluations of answers from GPT-4, LLaMA 3, Gemini, and online human therapists on patient questions from the public forum CounselChat. Each answer is rated across six clinically grounded dimensions, with span-level annotations and written rationales. Expert evaluations show that while LLMs achieve high scores on several dimensions, they also exhibit recurring issues, including unconstructive feedback, overgeneralization, and limited personalization or relevance. Responses were frequently flagged for safety risks, most notably unauthorized medical advice. Follow-up experiments show that LLM judges systematically overrate model responses and overlook safety concerns identified by human experts. To probe failure modes more directly, we construct CounselBench-Adv, an adversarial dataset of 120 expert-authored mental health questions designed to trigger specific model issues. Expert evaluation of 1,080 responses from nine LLMs reveals consistent, model-specific failure patterns. Together, CounselBench establishes a clinically grounded framework for benchmarking LLMs in mental health QA.
[Uncaptioned image] Code & Data: https://github.com/llm-eval-mental-health/CounselBench
Disclaimer: this study is conducted for research purposes and is not intended for clinical use or as a substitute for professional medical advice.
1 Introduction
Large language models (LLMs) have shown impressive progress in general-purpose question answering (QA), and there is growing interest in their use for healthcare applications (Thirunavukarasu et al., 2023). However, deployment in high-stakes settings requires rigorous evaluation of both capability and risk (Ali et al., 2023). Existing medical QA benchmarks such as MedQA (Jin et al., 2020) and MedMCQA (Pal et al., 2022) largely emphasize multiple-choice or fact-based tasks (Zhang et al., 2025). While these benchmarks measure factual recall, they fail to capture how models perform on the open-ended questions that patients actually ask. Open-ended QA presents greater challenges: questions are ambiguous, responses cannot be reduced to a single correct answer, and evaluation requires criteria that extend beyond factual accuracy.
These challenges are particularly critical in mental health QA. Here, patient queries are expressed as free-text descriptions of symptoms, treatment concerns, and emotional needs, and helpful responses must balance empathy, actionable guidance, and professional boundaries (Elliott et al., 2011; Na, 2024; Meadi et al., 2025). Unlike other medical domains, where diagnostic tests or structured knowledge often constrain the answer space, mental health QAs are inherently subjective and context-dependent (Association, 2013; Hormazábal-Salgado et al., 2024). Moreover, mental health QA has direct real-world significance: platforms such as CounselChat (Bertagnolli, 2020), peer-support forums (Pendse et al., 2019), EHR-based messaging, and digital mental health services like NOCD (NOCD, 2018) primarily operate through single-turn interactions, with limited follow-up in most cases. These setting are designed for brief, asynchronous exchanges that offer emotional support, psychoeducation, or coping strategies in a concise format. Such platforms are widely used and often serve as a primary source of support, especially for individuals who may not have access to or feel ready for full-length therapy (Naslund et al., 2016; Merchant et al., 2022). In such settings, even small errors, such as speculative claims, unsafe recommendations, or dismissive tone, can have immediate negative consequences. Evaluating LLMs on authentic, open-ended mental health QA in a standardized framework that enables consistent comparison of model responses to identical patient questions is therefore both technically and socially urgent.
While recent work has begun to explore mental health QA, current resources remain limited. Many benchmarks continue to rely on multiple-choice proxies, which bypass the ambiguity and contextual demands of free-text responses (Racha et al., 2025). Others use small expert panels or LLM-as-judge protocols (Sun et al., 2021; Szymanski et al., 2024; Racha et al., 2025), approaches that provide efficiency but often miss clinically salient failures. These limitations highlight the need for a large-scale, clinician-grounded benchmark for open-ended mental health QA, as well as resources designed to stress-test LLMs and surface failure modes that matter in practice.
We introduce CounselBench to address this gap. Our contributions are threefold:
  1. 1.
    Clinically grounded evaluation dimensions: we define six criteria for open-ended mental health QA, capturing both quality and safety.
  2. 2.
    CounselBench-Eval: we recruit 100 licensed or professionally trained mental health professionals to provide 2,000 evaluations of answers to real patient questions. These evaluations cover 4 answer providers (GPT-4, LLaMA-3, Gemini, and online human therapists) across 100 questions, with 5 independent expert annotations per response. Each annotation includes span-level labels and written rationales.
  3. 3.
    CounselBench-Adv: we construct an adversarial dataset of 120 expert-authored mental health questions designed to elicit systematic failure modes. For these questions, we collect 1,080 responses from nine models and have them evaluated by professionals, which identifies targeted signals for model improvement.
Together, these contributions move evaluation beyond factual correctness toward clinically meaningful behavior, establishing a practitioner-anchored benchmark for LLMs in mental health QA.
2 Related Work
LLMs in Medical QA Benchmarks. Early medical QA benchmarks emphasized closed-form accuracy, such as multiple-choice exams (Pal et al., 2022) or retrieval-based tasks (Zhu et al., 2020). Recent benchmarks broaden evaluation beyond accuracy with multi-axis rubrics. For example, MultiMedQA (Singhal et al., 2022) introduces physician-designed criteria such as factuality and bias. HealthBench (Arora et al., 2025) further scales to tens of thousands of physician-curated items in general medical domains. However, they remain focused on structured medical knowledge or rely on complex schemes that are difficult to apply in practice. In contrast, our work targets authentic mental health QA, where patient-facing questions require safety, empathy, and contextual sensitivity.
Evaluating Open-Ended Medical QA. Evaluating free-form medical QA remains difficult. Automated metrics (e.g., BLEU) scale easily but poorly capture qualities like specificity or safety (Croxford et al., 2024). Some studies incorporate expert review, but often with small panels due to cost (Hosseini et al., 2024; Kim et al., 2024; Manes et al., 2024). LLM-as-judge approaches are emerging (Szymanski et al., 2024), yet their reliability in healthcare domains requires human validation.
Benchmarks for Mental Health QA. Existing mental health QA datasets mostly focus on exam-style knowledge tests with objective answer keys (Racha et al., 2025; Nguyen et al., 2025). Others adopt psychological rubrics (Li et al., 2024), but rarely involve clinicians in design (Chiu et al., 2024); the evaluation is often scaled by non-expert raters or heuristics (Young et al., 2024). Our evaluation protocol is co-designed with clinical psychologists and scaled annotations to 100 licensed or trained professionals, ensuring clinical validity and reproducibility.
Adversarial Evaluation in Mental Health QA. Prior studies have highlighted risks of LLMs in mental health support, such as unsafe advice or discrepant responses based on inferred demographics (Lawrence et al., 2024; Imran et al., 2023; Gabriel et al., 2024). However, most findings come from post hoc analyses of deployed systems (Song et al., 2025; Li et al., 2025b), which reveal problems retrospectively rather than prospectively stress-testing models. A few recent works adapt red-teaming to mental health (Grabb et al., 2024; Schoene & Canca, 2025), but their failure modes are predefined and literature-driven, limiting coverage of the diverse failure patterns that emerge in practice. In contrast, our adversarial dataset is empirically grounded: 10 mental health professionals authored 120 prompts from failures observed in CounselBench-Eval, creating an expert-crafted benchmark for prospectively eliciting clinically meaningful vulnerabilities.
3 Overview of CounselBench
We construct CounselBench in collaboration with 100 mental health professionals. It is a two-part benchmark for evaluating LLMs in realistic, mental health QA scenarios (Figure 1):
  • COUNSELBENCH-Eval: a benchmark of 2,000 expert evaluations of LLMs and online human therapist responses to real patient queries, with numeric scores, span-level annotations, and rationales.
  • COUNSELBENCH-ADV: an adversarial dataset of 120 expert-authored questions designed to expose identified failure modes in LLMs, paired with 1,080 responses from 9 LLMs, each annotated for whether the targeted failure mode is present.
Figure 1: Overview of CounselBench benchmark. COUNSELBENCH-EVAL (left) includes expert evaluation of LLMs and online human therapist responses to real counseling questions. COUNSELBENCH-ADV (right) includes adversarial questions authored by clinicians to target identified LLM failure modes. See Appendix B for license/degree types and specialization areas.
Research Pipeline. We began by selecting 100 real patient questions from the public forum CounselChat (Bertagnolli, 2020), spanning 20 common mental health topics (§4.1). Each question was answered by GPT-4, Gemini, LLaMA-3, and an online human therapist (§4.2). To evaluate responses, we defined six evidence-based dimensions grounded in literature: overall quality, empathy, specificity, factual consistency, medical advice, and toxicity (§4.3). We then recruited 100 mental health professionals for evaluation (§4.4). Besides assigning ratings, annotators provided span-level annotations and written rationales, yielding rich insights into response quality and safety. These evaluations form the COUNSELBENCH-EVAL dataset. We also prompted nine LLMs to rate the same responses under the same rubric (§4.6), enabling systematic comparison between human and LLM judges. To probe safety directly, 10 experts authored 120 adversarial questions and additional 5 expert annotated their identified failure modes, creating the CounselBench-Adv dataset (§5), to trigger failure patterns identified in CounselBench-Eval.
4 CounselBench-EVAL: Expert Evaluation of LLMs for Open‑Ended Mental‑Health QA
We develop COUNSELBENCH-EVAL to systematically assess LLM behavior on authentic mental health questions. This section describes the construction pipeline: real-world mental health QA sourcing (§4.1), response generation (§4.2), rubric design (§4.3), human annotation protocol (§4.4), results from expert scoring and analysis (§4.5), and experiments with LLM judges (§4.6).
4.1 Real-World Mental Health Question and Answering
We source mental health QA pairs from CounselChat, a platform where licensed therapists publicly respond to anonymous patient questions (Bertagnolli, 2020). All posts were authored before 2022, ensuring that all online therapist responses predate the widespread use of LLMs like ChatGPT.
Question Curation and Topic Coverage. To ensure both topical diversity and response quality, we selected the topics and questions based on predefined criteria (Appendix  LABEL:appendix_question_filtering). From each topic, we selected five questions with the highest response upvotes, resulting in 100 questions in total. The topics were: (1) depression, (2) relationships, (3) anxiety, (4) family conflict, (5) parenting, (6) self esteem, (7) relationship dissolution, (8) behavioral change, (9) anger management, (10) trauma, (11) marriage, (12) domestic violence, (13) grief and loss, (14) social relationships, (15) workplace relationships, (16) legal regulatory, (17) substance abuse, (18) counseling fundamentals, (19) eating disorders, (20) professional ethics. The keywords for each question are summarized in Table LABEL:tab:keyword_per_question.
Human Answers. All responses on CounselChat are written by verified, non-anonymous therapists with public profiles; only patient questions are anonymous (Bertagnolli, 2020). Despite that, forum contributions are informal and vary in quality. To reduce this variability and focus on consistently well-regarded content, we selected the top-voted answer for each question (Appendix  LABEL:appendix_human_responses). Upvotes are a standard, publicly observable indicator of perceived helpfulness and relevance in peer-support forums, and have been widely used in prior work to capture perceived helpfulness(Jiang et al., 2020; Huang et al., 2025), social support (De Choudhury & De, 2014), engagement (Gao et al., 2020), community reception (Alghamdi et al., 2025), and narrative quality (Wang et al., 2017).
4.2 Generating Counseling Responses from LLMs
Pilot Study. We first piloted domain-specific models such as MentalLLaMA (Yang et al., 2023) and Meditron (Chen et al., 2023b). As shown in Appendix LABEL:sec:domain-specific-experiments, these models performed poorly on open-ended mental health QA, likely due to narrow fine-tuning or lack of instruction tuning. We therefore focused on general-purpose LLMs that represent current deployment paradigms .
Model Selection. To ensure fair comparison and reduce annotator fatigue, we limited our evaluation to three LLMs per question. Each mental health professional blindly annotated four answers (three LLM-generated and one human-written), enabling consistent evaluation while keeping the task cognitively manageable. The selected models are: (1) GPT-4-0613 (via OpenAI API), a widely adopted commercial model, (2) LLaMA-3.3-70B-Instruct (via Hugging Face), a leading open-source, instruction-tuned model from Meta, (3) Gemini-1.5-Pro (via Google Generative AI SDK), which powers responses in Google Search and Assistant. Together, they represent major deployment paradigms: commercial APIs, open-source checkpoints, and search-integrated platforms. Each model is prompted with the original patient question and asked to respond. Full prompt templates and generation details are provided in Appendix D.1.
4.3 Evaluation Rubric and Paradigms
We developed a multi-dimensional evaluation rubric based on clinical psychology literature and expert consultation. All metrics were rated using 5-point Likert scales ($1=$ the most negative; $5=$ the most positive) unless specified. Appendix A shows the rating interface.
Overall Quality captures the holistic judgment of each response.
Empathy measures whether the response demonstrates emotional attunement, compassion, or validation. Empathy is a central component of therapeutic modalities (e.g., person-centered therapy (Jani et al., 2012)), and is strongly linked to therapeutic alliance, a key predictor of positive therapy outcomes (Ackerman & Hilsenroth, 2003).
Specificity evaluates how well the response is tailored to the user’s particular context, rather than offering generic or overly broad advice. Personalized replies help users feel heard and understood (Elkin et al., 2014; Stiles & Horvath, 2017; Kramer & Stiles, 2015), and contextual relevance is critical to effective support (Hitchcock et al., 2024; Stiles & Horvath, 2017; Kramer & Stiles, 2015).
Medical Advice flags whether the response includes therapeutic or diagnostic guidance that should only be provided by licensed professionals. Such advice typically requires individualized assessment and clinical expertise (Frank et al., 2020; Lyon et al., 2011), and its timing and appropriateness must be carefully managed by trained practitioners (Hill et al., 2023; Prass et al., 2021). This dimension is treated as binary (Yes/No) with an additional “I am not sure” option because fine-grained distinctions in severity and subcategories are difficult to assess consistently for Medical Advice. We further ask annotators to provide the exact advice spans and written rationales, which enable post hoc analysis of different categories, and levels of severity. Higher rates of flagged content indicate safety concerns.
Factual Consistency checks whether the response aligns with accepted commonsense or clinical knowledge and avoids inaccurate or unsupported claims. Users often struggle to discern factual accuracy in online mental health content (Morahan-Martin & Anderson, 2000). LLMs are known to hallucinate false information (Sun et al., 2025), and even human counselors can produce factual errors in communication (Lobel, 2024). This dimension is rated on a 4-point scale, also with an additional “I’m not sure” option to accommodate cases where the rater cannot confidently assess it.
Toxicity assesses the presence of language that is potentially harmful, stigmatizing, dismissive, or ethically problematic. Individuals seeking mental health support are often in vulnerable states and may be particularly susceptible to damaging content (Webb et al., 2008; Mitchell et al., 2017).
Rating Rationale & Text Extraction To capture the context and reasoning behind human evaluations of mental health QA, we asked annotators (§4.4) to provide open-text explanations for their ratings on Overall Quality, Medical Advice, Factual Consistency, and Toxicity. Additionally, we asked them to identify and extract specific portions of each response that supported their ratings. Two researchers on the team checked and verified the quality of text responses after data collection. In general, human annotators provided detailed written feedback, with an median length of 576.5 written words (IQR: 339.8 - 866.5) per survey combining all rationales.
Figure 2: Distribution of (A) credential types and (B) counseling experience among the 100 annotators.
4.4 Expert Annotation Protocol
We implemented the evaluation rubric as a structured online questionnaire using Qualtrics, a widely used platform for human subjects research.
Pilot Studies. Prior to full deployment, we validated our questionnaire and study protocol with 8 participants through three rounds of pilot studies. Accordingly, we refined our evaluation questions and improved their clarity. For example, some experts occasionally expressed uncertainty on Medical Advice and Factual Consistency, which is not observed in other dimensions, so we included the option ”I am not sure” in the main study to reduce noise and avoid forcing uncertain judgments.
Recruitment and Credential Verifications. We recruited 100 licensed or professionally trained mental health practitioners across the U.S. through Upwork, the world’s largest talent platform commonly used in human-computer interaction (HCI) research for sourcing domain-expert annotators (Fulker & Riedl, 2024). All annotators had formal counseling training or certification, and we individually verified their education, credentials, or professional experience (Appendix B).
Annotation Protocol. Each annotator was assigned a randomly sampled survey containing 5 counseling questions, each paired with 4 responses: one from an online human therapist and three from different LLMs. The four responses for each question are in randomized order to mitigate position bias. This resulted in 20 QA pairs per annotator. Each pair was independently evaluated by five different professionals, supporting inter-rater reliability and reducing individual bias. For each question, all four responses were rated by the same group of annotators, enabling direct and fair comparison across models. In total, the study includes 2,000 annotated QA pairs (100 questions × 4 answers × 5 annotations). Annotators were blind to the source of each response and were not informed that some had been generated by LLMs.
Annotator Demographics and Engagement. Collectively, the 100 annotators held 32 distinct types of licenses and degrees, and reported expertise spanning 43 specialized counseling areas (summarized in Figure 2; more details in Appendix B). This broad coverage ensures that our evaluation reflects a diverse and representative cross-section of real-world practice. The demographic composition of our annotator sample aligns with national statistics for the U.S. counseling profession, which shows a higher proportion of white and female practitioners (McGuire & Miranda, 2008; Verdieu, 2024; American Psychological Association, 2022). The median written rationale length was 576.5 words across 20 QA pairs (IQR: 339.8 - 866.5), and the median time spent was 1 hr 22 min (IQR: 52 min - 3 hrs 9 min), indicating substantial engagement.
4.5 Expert Evaluation Results: Counseling Quality Across LLMs and Online Human Therapists
We present the results of expert annotations comparing the responses generated by GPT-4, LLaMA-3.3, Gemini-1.5-Pro, and online human therapists.
Overall Performance. Table 1 reports the average expert-annotated scores for each model. Among the models, LLaMA-3.3 received the highest overall ratings, leading on five of six dimensions. Pairwise significance tests (Woolson, 2005) are reported in Appendix LABEL:appendix_pairwise_significance_test. Despite strong scores, a higher proportion of LLaMA-3.3’s outputs (14%) were flagged for providing unauthorized medical advice (e.g., recommending therapy techniques), highlighting an important safety concern. In contrast, GPT-4 responses were more likely to include safety disclaimers, with approximately one-third of outputs explicitly declining to answer and recommending consultation with human professionals.
Table 1: Average expert ratings of counseling responses across six evaluation criteria. Each response was rated by five mental health professionals; scores were first averaged by question, then by model. Responses marked ”I am not sure” were excluded. For Medical Advice, the percentage of “Yes” responses was computed per question (excluding “I’m not sure”) and averaged over all questions.
Overall $\uparrow$ (range:1-5) Empathy $\uparrow$ (range: 1-5) Specificity $\uparrow$ (range: 1-5) Medical Advice (percentage of “Yes”) Factual Consistency $\uparrow$ (range: 1-4) Toxicity $\downarrow$ (range: 1-5) GPT4     $3.28$     $3.37$     $3.46$      0.07      $3.53$     $1.78$ Llama 3.3     $4.29$     $4.22$     $4.63$      0.14      $3.70$     $1.36$ Gemini-1.5-Pro     $3.26$     $2.76$     $3.50$      0.08      $3.52$     $1.64$ Online Human Therapists     2.60     2.72     3.29      0.17      2.92     2.56
Table 2: Mean (over questions) Krippendorff’s alpha (ordinal); not applied to binary Medical Advice.
Overall Score Empathy Specificity Factual Consistency Toxicity Krippendorff’s alpha (K-$\alpha$) 0.82 0.83 0.82 0.75 0.72
Annotation Consistency. We verified the consistency of our expert annotators’ ratings by calculating inter-rater reliability (Table 2); annotator responses across measured metrics yielded Krippendorff’s alpha values  $\geq 0.7$, indicating a high level of inter-rater agreement (Zapf et al., 2016).
Annotation Analysis. Additional analyses of annotator behavior and score distributions are reported in Appendix C. We find that annotators who spent more time (top 50%) wrote longer rationales and those in the top 25% of time spent also labeled more sentences as Medical Advice. Significant differences in Empathy, Medical Advice, and Factual Consistency were also observed across annotators with different years of experience. Finally, Table LABEL:tab:uncertain-percentage reports the fraction of “I am not sure” selections, which were infrequent overall.
Qualitative Analysis of Annotator Rationales. To better understand expert judgments, we conducted a thematic analysis following established protocol (Clarke & Braun, 2017). Two researchers on our team independently reviewed and thematically coded annotators’ open-text explanations. Seven primary issues among low-quality responses are identified (listed in Appendix LABEL:sub_appendeix_llm_judge_eval). We then prompted GPT-4.1 to label all low-rated responses (overall score $\leq$ 2) using these categories. Each response could exhibit multiple issues. A detailed breakdown is provided in Figure LABEL:fig:low_overall (Appendix LABEL:sub_appendeix_llm_judge_eval). The top two most frequently flagged failure types for each model are:
  • GPT-4: Offering unconstructive feedback (e.g., lacking clarity or actionability) (49.3%) and demonstrating little personalization or relevance (41.3%).
  • Llama-3.3: Overgeneralizing or making judgments and assumptions without sufficient context (66.7%) and offering unconstructive feedback (e.g., lacking clarity or actionability) (27.8%).
  • Gemini-1.5-Pro: Lacking empathy or emotional attunement (44.1%) and offering unconstructive feedback (e.g., lacking clarity or actionability) (40.7%).
  • Online human therapist: Overgeneralizing or making judgments and assumptions without sufficient context (46.7%) and displaying an inappropriate tone or attitude (e.g., dismissive, superficial) (36.1%).
Clinical Boundaries: Unauthorized Medical Advice in LLM responses. To further examine safety risks, we analyzed the response excerpts flagged by expert annotators for containing medical advice. The most concerning identified issue was the inclusion of unauthorized clinical recommendations, especially prescribed psychotropic treatments (i.e., medications for mental health treatments, according to National Institute of Mental Health (of Mental Health, 2025)).
Among the LLMs, GPT-4 occasionally recommended specific medications, such as selective serotonin reuptake inhibitors (SSRIs). Annotators also identified that all three LLMs occasionally advised therapy techniques like cognitive behavioral therapy (CBT) or mindfulness. While these practices are common in clinical care, professionals emphasized that they require tailored application and oversight. In CounselBench-ADV, we further refined this category into (i) recommending specific medications, (ii) suggesting therapy techniques, and (iii) speculating about medical symptoms.
4.6 Can LLMs Reliably Judge the Quality of Responses?
Human expert evaluation provides gold-standard insight but is costly and difficult to scale (Gehrmann et al., 2021). As LLMs are increasingly considered for deployment in mental health applications, one important question is whether models can serve as their own evaluators. Prior work has proposed the “LLM-as-Judge” paradigm for tasks such as summarization and factuality (Madaan et al., 2023; Manakul et al., 2023), but its reliability in high-stakes, subjective domains like mental health remains unclear. To investigate this, we tested nine advanced LLMs as automated judges, as listed in Figure 3. Each model was prompted with the same QA pairs used in Section 4.4, along with the evaluation criteria provided to human experts. Prompts followed the survey format used in our human annotation protocol, with minor formatting constraints tailored to each model (see Appendix LABEL:sub_appendeix_llm_judge_eval), and additional results for LLM judges are in Appendix LABEL:llm_judge_additional.
Figure 3: Average evaluation scores across six dimensions (subplots) for responses generated by GPT-4, LLaMA-3.3, Gemini-1.5-Pro, and online human therapists (x-axis in each subplot). Each colored line represents one evaluator, including nine LLM-based judges and human experts (red). Higher values indicate better performance except for Toxicity and Medical Advice. See Table LABEL:llm_judge_numeric_score for full numerical results.
LLM Judges Overrate Responses and Miss Safety Failures. Figure 3 presents the scores provided by LLM judges, compared with human expert annotations. Most LLM-based evaluations yielded inflated scores relative to human ratings. This was most pronounced for Factual Consistency, where LLM judges frequently assigned perfect or near-perfect scores to LLM-generated responses, regardless of their actual content. Ratings for Toxicity showed minimal variance across all sources. All LLM judges uniformly assigned the lowest toxicity scores, even when human experts had flagged content as potentially harmful or inappropriate. This lack of sensitivity indicates a critical weakness in the use of LLMs for safety evaluations in high-risk domains.
LLM Rankings Diverge from Human Preferences. Beyond absolute scores, we compared model rankings based on average quality, as shown in Figure 3. The preferences expressed by LLM judges diverged sharply from those of human annotators, except for GPT-5 (dark blue line in Table 3), which calls into question the reliability of LLMs in identifying relative quality. Most notably, Gemini-1.5-Pro was rated as the lowest-performing model by human expert annotators but was consistently ranked above GPT-4 by every LLM judge for Overall Quality.
LLM Judges Rarely Flag Problematic Text. We also tested whether LLM judges could identify problematic content at the sentence level. As in the human evaluation protocol, each model was asked to extract spans containing medical advice, factual errors, or toxic language. We compared the extracted text against sentences flagged by at least two human annotators, treating these as widely agreed indicators of failure (more details in Appendix LABEL:sec:sentence_level). Table LABEL:tab:miss_percentage in the Appendix reports the proportion of widely-flagged sentences detected by each LLM judge. Most models failed to catch any toxic or incorrect sentences. Some caught limited instances of unauthorized medical advice, but none showed robust or reliable detection across safety categories. These results help explain the inflated consistency and toxicity scores and highlight the limitations of LLM-based judging.
5 CounselBench-ADV: An Adversarial Benchmark for Surfacing LLM Failures
While expert annotators rated LLM-generated responses to open-ended mental health questions as generally high quality, our in-depth analysis revealed recurring issues, such as inappropriate advice or generic feedback (§4.5). To probe these failure modes more systematically, we introduce CounselBench-ADV, an adversarial benchmark constructed by mental health professionals.
Identifying Fine-grained Failure Modes. To construct a targeted adversarial benchmark, we first need to isolate the failure modes that LLMs exhibit. Broad categories (e.g., “lacking personalization”) from Section 4.5 are too general. To enable precise probing, we conducted a deeper review of expert rationales and flagged responses to extract concrete failure modes:
  • GPT-4: provides specific medication (1. medication) and suggests specific therapy techniques (2. therapy).
  • Llama-3.3: speculates about medical symptoms (3. symptoms) and is judgmental (4. judmental).
  • Gemini-1.5-Pro: is apathetic (5. apathetic) and is based on unsupported assumptions (6. assumptions).
Here we selected two per model that best represent the kinds of specific issues each model (see Table LABEL:tab:six_issues in Appendix LABEL:appendix_failure_modes). Each issue is linked to real examples from CounselBench-EVAL, providing clear foundations for clinicians to construct new questions in a controlled and measurable way.
Adversarial Dataset Construction. We rehired 10 mental health professionals who previously participated in CounselBench-EVAL. Each clinician was provided with detailed definitions and real examples illustrating each failure type, drawn directly from flagged responses in CounselBench-EVAL (Table LABEL:tab:six_issues). They were instructed to independently author two realistic mental health questions for each failure mode, i.e., crafted in such a way that it would be plausible for an LLM to fall into the corresponding error (see the questionnaire screenshot in Appendix 4). Each clinician authored 12 adversarial questions (2 per issue × 6 issues), resulting in a total of 120 questions. Clinicians were asked to ensure the realism and diversity of the questions, reflecting what actual patient might ask. Importantly, these questions do not directly contain the failure, but are designed to trigger the failure in LLM responses. The resulting dataset provides a high-precision probe of specific vulnerabilities in LLM-generated responses.
Evaluation Setup. We used the same prompting setup as in CounselBench-EVAL for these adversarial questions with nine LLMs. For each model, we generated one response for each question, this leads to 1,080 QA pairs (120 questions × 9 LLMs). The annotation task was deliberately framed as classification (“yes”/ “no”/“not sure”) with clear clinician-authored definitions and one in-context example per failure type (Appendix LABEL:appendix_failure_modes), rather than the multi-dimensional metrics used in CounselBench-Eval. Five mental-health practitioners, different from those who authored the adversarial questions, then annotated whether each response has the targeted issues. The full annotation setup is described in Appendix LABEL:sec:adv_llm_judge_results.
Table 3: Fraction of responses identified to contain each targeted failure mode by five mental health professionals. Higher values reflect greater model vulnerability to the targeted issue.
Specific Issue GPT-3.5-Turbo GPT-4 GPT-5 Llama-3.1 Llama-3.3 Claude-3.5-Sonnet Claude-3.7-Sonnet Gemini-1.5-Pro Gemini-2.0-Flash 1. Medication   0.05 0   0.47   0.05   0.1 0 0 0 0 2. Therapy   0.2   0.2   0.85   0.55   0.65   0.45   0.5   0.2   0.26 3. Symptoms   0.15   0.45   0.6   0.45   0.45   0.5   0.37   0.26   0.25 4. Judgmental   0.25   0.25   0.05   0.11   0.1   0.05   0.1   0.2   0.1 5. Apathetic   0.7   0.2   0.15   0.15   0.15   0.05   0.2   0.4   0.3 6. Assumptions   0.4   0.35   0.15   0.25   0.25   0.35   0.25   0.4   0.35
Adversarial Questions Effectively Trigger Targeted Failures. Table 3 reports the frequency of each failure mode across model outputs based on expert evaluation. The adversarial questions effectively elicited the intended issues in a large fraction of cases. Several patterns stand out: therapy suggestions are highest for GPT‑5 (0.85), substantial for Llama models (0.55–0.65), and present for Claude models (0.45–0.50). Symptoms speculation is also frequent in the three above model families, and lower for Gemini (0.25-0.26) and GPT‑3.5‑Turbo (0.15). Unsupported assumptions appear in most models (0.25–0.40), with GPT‑5 lower at 0.15. Medication advice is rare for nearly all models (0–0.10), but GPT‑5 is an outlier at 0.47. Apathetic tone peaks for GPT‑3.5‑Turbo (0.70), while judgmental tone remains comparatively low across models (0.05–0.25).
Interestingly, the failure modes revealed by CounselBench-ADV are not randomly distributed across models, but exhibit clear family-level patterns. Models within the same family (e.g., LLaMA, Gemini, and Claude) have similar distributions across the failure modes, while GPT family show distinct patterns. This suggests that failure modes remain relatively consistent within model families, but can shift substantially across major model upgrades (OpenAI et al., 2024; Chen et al., 2023a).
Table 4: LLM-as-Judge performance on CounselBench-Adv, relative to human expert ratings.
Metrics LM-as-Judges GPT-3.5-turbo GPT-4 GPT-4.1 GPT-5 Llama-3.1-70B-Instruct Llama-3.3-70B-Instruct Claude-3.7-sonnet Gemini-2.0-flash Acc. 0.74 0.70 0.64 0.67 0.63 0.64 0.70 0.63 F1 0.41 0.35 0.49 0.49 0.48 0.48 0.50 0.46
LLM Judges Remain Misaligned with Human Expert Judgments.
We also evaluated an LLM-as-Judge setup for failure-mode detection across eight models  [^1] Gemini-1.5-Pro and Claude-3.5-Sonnet, which are included in the LLM-as-judge analysis in Section §4.6, were not longer available when we ran the LLM-as-judge experiments on CounselBench-ADV. . The experiment follows the same setup in Section §LABEL:sub_appendeix_llm_judge_eval with LABEL:prompt:gpt4.1-failure-mode, and additional details are provided in Appendix LABEL:sec:adv_llm_judge_results. Despite providing explicit definitions and in-context examples, we observe a substantial gap between LLMs and expert judgements (Table 4). Even the best-performing model (Claude-3.7-Sonnet) attains an F1 score of only 0.5. The performance breakdown of each LLM judge for each failure mode is present in Figure LABEL:fig:tpr_fpr_llm_judge_adv. Additional analysis for the LLM judges is provided in Appendix LABEL:sec:adv_llm_judge_results.
6 Discussion
This work presents CounselBench, a benchmark that evaluates LLMs on open-ended mental health QA. Unlike prior resources that rely on multiple-choice tasks or small-scale expert review, it brings together two complementary components: a large-scale corpus of 2,000 expert evaluations with detailed written rationales (CounselBench-Eval), and a set of 120 adversarial questions written by clinicians to expose failure modes paired with 1,080 expert annotations for failure mode detection (CounselBench-Adv). Together, they provide both breadth and depth: broad coverage through systematic expert scoring, and targeted probing through adversarial design. Our analysis shows that LLMs can achieve high ratings on several dimensions of quality, yet continue to exhibit recurring weaknesses, including unconstructive feedback, overgeneralization, limited personalization, and unauthorized medical advice. These findings highlight both the progress and the open challenges in deploying LLMs responsibly in mental health QA.
As noted in prior surveys of mental‑health datasets (Harrigian et al., 2021; Mandal et al., 2025), publicly available data with clinician responses are extremely scarce. On one hand, most clinical conversations and patient-provider messages cannot be released due to privacy protections (e.g., de-identification risks highlighted by Kovačević et al. (2024)); on the other hand, peer‑support forums (Marshall et al., 2024) and social‑media platforms such as Reddit (Boettcher, 2021) typically contain non‑clinical, user‑generated content with noisy labels and without professional verification (Garg, 2023). Within this landscape, CounselChat remains one of the few openly accessible resources that pair patient questions with verified therapist responses, and prior work has examined its relatively high quality and utility for research (Ahmed et al., 2024). Although current public datasets are limited, the evaluation framework in CounselBench, including the six-dimension rubric and annotation protocol, can be applied directly to future datasets. This structure can support consistent model comparison across broader settings in subsequent work.
While CounselBench is designed for evaluating open-ended mental health QA, our framework provides a foundation for future expansion into dialogue settings. For CounselBench-Eval, the rubric can be applied turn-by-turn or at the session level (Chen et al., 2025) and aggregated across the dialogue, for example, by averaging scores or weighting them based on conversational role (e.g., first vs. follow-up turns). Moving to multi-turn exchanges, however, introduces additional complexity, including the need for models to appropriately track and utilize prior conversational context. This raises challenges in both prompting design (e.g., how much history to include) and evaluation (e.g., how to measure coherence or consistency over time). For CounselBench-ADV, extending to multi-turn adversarial prompts requires preserving interactional dynamics while targeting specific failure modes. Promising directions include using simulated patient agents (Wang et al., 2024a; b) to scaffold coherent multi-turn interactions that elicit subtle failures, and developing multi-turn red-teaming techniques seeded from our identified adversarial examples (Guo et al., 2025).
Both components of CounselBench are released as public resources to support future research. CounselBench-Eval can be used to (1) train and validate alignment methods such as reinforcement learning from human feedback (Ouyang et al., 2022; Zhang et al., 2024; Li et al., 2025a), (2) develop automated quality or safety detectors grounded in span-level annotations (Zhang et al., 2023), and (3) study expert rationales to build critique models or interpretability tools. CounselBench-Adv enables complementary research, such as (1) stress-testing new models to compare robustness under clinically relevant failure modes, (2) incorporating adversarial cases into red-teaming and auditing pipelines (Mazeika et al., 2024), (3) using these targeted examples as hard negatives for fine-tuning or as feedback signals in multi-agent critique systems (Paulus et al., 2025; Gomaa et al., 2025). By releasing these resources, we aim to support both rigorous benchmarking and practical model improvement in mental health QA and related high-stakes applications.
Ethics Statement
This study was reviewed and approved as exempt research by the Institutional Review Board of University of Southern California (Study ID: UP-25-00022). Questions and clinician responses are sampled from MIT-licensed CounselChat dataset (Bertagnolli, 2020). All identifiers (e.g., names, homepage links) were removed prior to use. Human annotators provided informed consent, including consent for release of de-identified data, and no personally identifiable information was collected.
Acknowledgments
We thank the mental health professionals who contributed to this project. Their careful annotations, clinical insights, and generous support were invaluable to our study. We also thank Athan Li, Konghao Zhao, and Shupeng Cheng for participating in the initial survey rounds and offering helpful input, and Maya Andres for reviewing and providing feedback on our early survey materials. Finally, we thank Bingsheng Yao and Dakuo Wang for their helpful discussions and feedback on COUNSELBENCH-EVAL.
Reproducibility statement
We release the data and source code. Full hyperparameters, implementation details, random seeds, and prompt templates are in Appendix D, and computational resources are detailed in Appendix LABEL:compute_resouces, to ensure replicability.
Appendix A Survey Design
We utilized Qualtrics to conduct our surveys. We built a template as shown in Figure 4 for CounelBench-Eval and then populated surveys for all question and response pairs. Each annotator evaluated five randomly sampled questions, each accompanied by four answers - one from GPT-4, Llama-3.3, Gemini-1.5-Pro, and an online human therapist. For each post-response pair, annotators evaluated across our metrics, providing rationals and span-level annotations. At the end of the survey, we collected demographic and professional background data, including gender, ethnicity, freelance capacity, counseling experience, credentials, and areas of specialization.
Figure 5 shows a screenshot for collecting CounselBench-ADV questions. Each annotator was asked to write two new questions per failure mode, which could trigger the targeted failure mode. The failure mode descriptions/examples/details for the questionnaires are listed in TableLABEL:tab:six_issues.
Figure 4: Survey interface CounselBench-Eval: annotators read a user post and one response (left) and rate the response on criteria with Likert scales and text‐evidence boxes (right).
Figure 5: Survey Interface for CounselBench-ADV: annotators are given the failure mode, examples, and explanations for the targeted failure mode and are asked to draft new questions that could trigger the same failure mode.
Appendix B Annotator Demographic Overview
We asked all applicants to provide when they apply: 1) List at least one relevant counseling certifications, licenses, or degrees. 2) Provide a brief overview of their counseling experience.
We further verified the credentials of our 100 expert annotators through multiple (disjoint) methods: (1) license numbers if provided (34 cases), (2) professional profile on platforms such as Psychology Today (38 cases), (3) working and education history on Linkedin (21 cases), (4) listings in workplace or school directories (3 cases), (5) employment and education history on public Upwork profiles (4 cases). We ensured that all annotators had completed counseling-related training or possessed relevant counseling experience.
Table LABEL:tab:license_degree_summary reports the distribution of counseling licenses and degree types, while Table 5 lists the specialization areas of the 100 mental-health professionals. Figure 6 shows the ethnicity and gender distribution of annotators. All data are from background questions in the survey.
Figure 6: Ethnicity (left) and gender distribution (right) of the 100 mental-health professional annotators.
Table 5: Distribution by Counseling Specialization. Annotators may have multiple specializations and thus may be counted more than once in this table.
Counseling Specialization Count
Depression/anxiety   77
Trauma/PTSD   62
Child/adolescent therapy   40
Crisis intervention   33
Grief counseling   23
Substance use disorders   22
Couples/family therapy   21
Obsessive-Compulsive Disorder (OCD)   4
Attention-Deficit/Hyperactivity Disorder (ADHD)   3
LGBTQ + LGBTQIA+ issues   3
Prefer not to answer   2
Autism Spectrum Disorder (ASD)   2
Relationships   2
Life transitions   2
Self-Esteem   2
School counseling   2
Perinatal mental health   2
Sexual Disorders   1
Dissociation   1
Bipolar   1
Holistic wellness   1
Chronic illness/pain   1
Use substances issues   1
Narcissistic abuse   1
Codependency   1
Gender   1
Sexual orientation   1
Young adults Eating disorders   1
Women’s issues   1
Geropsychology   1
Neurodiversity   1
Exposure and Response Prevention(ERP)   1
Cognitive Behavioral Therapy(CBT)   1
Eating disorder/body dysmophia/HAES counseling   1
Religious trauma   1
Relational concerns   1
Identity development   1
Neuro-affirming mental health therapy   1
Disabilities   1
Corrections   1
Disordered eating   1
Other diverse identities presented in multicultural counseling   1
Psychosis   1
Mood disorders   1
Table 6: Annotator Degrees, Licenses, and Additional Certifications. Annotators may hold multiple credentials and may therefore appear more than once in the table.
Licensed Professionals Count
Licensed Professional Counselor (LPC)    25
Licensed Clinical Social Worker (LCSW)    18
Licensed Marriage and Family Therapist (LMFT)    8
Licensed Mental Health Counselor (LMHC)    5
Licensed Addiction Counselor (LAC)    2
Licensed Independent Social Worker of Clinical Practice (LISW-CP)    1
Licensed Clinical Professional Counselor (LCPC)    1
Clinical Supervisor (CS)    1
Licensed Addiction Counselor Supervisor (LACS)    1
Doctoral-Level Degrees
Psychologist (Ph.D./Psy.D.)    10
Psychiatrist (MD)    1
Trainees / Associates / Under Supervision
Licensed Associate Professional Counselor (LAPC)    3
Licensed Master Social Worker (LMSW)    2
Licensed Professional Counselor Associate (LPCA)    1
Licensed Mental Health Counselor Associate (LMHCA)    1
Licensed Associate Counselor (LAC)    1
Licensed Professional Counselor Candidate (LPCC)    1
Licensed Clinical Social Worker Associate (LCSWA)    1
Licensed Psychological Associate (LPA)    1
Licensed Social Worker (LSW)    1
Licensed Alcohol and Drug Counselor (LADC)    1
Currently in training/supervision   11
Other Degrees
Education Specialist Degree in school psychology    2
Master’s Degree in Counseling    1
Master’s Degree in School Counseling    1
Master’s Degree in Counseling Psychology    1
Master’s Degree in Psychology    1
B.A. in psych   1
Formerly Licensed / Associate
Licensed Associate Counselor (LAC)   1
Licensed Marriage and Family Therapist (LMFT)   1
Counseling-Related Certifications
National Certified Counselor (NCC)    2
School Psychologist License   1
K-12 School Counselor certification   1
Certified in FFT (Functional Family Therapy)   1
Trauma Informed CBT Certified   1
Appendix C Annotation Analysis
C.1 Annotation Time and Rationale Length
For quality evaluation, we recorded both the length of the written rationale and time spent per annotator. The median written rationale length was 576.5 words across 20 QA pairs (IQR: 339.75 - 866.5), and the median time spent was 1 hr 22 min (IQR: 52 min - 3 hrs 9 min), indicating substantial engagement.
To examine how time relates to scoring, we binned annotators into quartiles by total time spent (0 – 25%, 25 – 50%, 50 – 75%, 75 – 100%) and compared outcome scores across these groups. To ensure a fair comparison, we focused on the 460 Q&A pairs rated by annotators from all four quartiles. The Table 7 shows the mean score per dimension across quartiles, along with the Kruskal-Wallis $p$-values (for ordinal ratings) and chi-squared test (for categorical variables, i.e. Medical Advice and Factual Consistency because of the option “I am not sure”).
Table 7: Mean metric scores by annotator time quartile (0–25%, 25–50%, 50–75%, 75–100%) for the 460 Q&A pairs rated by annotators from all quartiles. The bottom row reports $p$-values for differences across quartiles (Kruskal–Wallis test for ordinal metrics; chi-squared test for categorical metrics).
Time Quartile Overall Empathy Specificity Medical Advice Factual Consistency Toxicity 0-25% (lowest) 3.29 2.93 3.58 0.13 3.48 1.82 25-50% 3.32 3.32 3.63 0.18 3.29 1.83 50-75% 3.36 3.11 3.77 0.07 3.38 1.83 75-100% (highest) 3.29 2.95 3.64 0.2 3.5 1.5 $p$-value 0.98 0.12 0.61 << .001 << .001 0.08
We found statistically significant associations for Medical Advice and Factual Consistency, which involve complex factual judgments. For example, annotators in the top quartile tended to give more conservative Medical Advice scores (Medical Advice=0.20) than the lowest quartile (Medical Advice=0.13). However, we emphasize that these are associational patterns, not causal effects. Longer time may indicate greater care or, alternatively, more difficult cases for the annotator.
Finally, we assessed whether time correlates with rationale length. The Table 8 shows the average word count per annotator across quartiles. As expected, annotators who spent more time tended to write longer rationales, suggesting greater annotation effort.
Table 8: Average total rationale word count per survey by annotator time quartile (0–25%, 25–50%, 50–75%, 75–100%). Differences across quartiles are significant (by the Kruskal–Wallis test); annotators who spent more time wrote longer rationales on average.
Time Quartile 0-25% (lowest) 25-50% 50-75% 75-100% (highest)
Average Word Count 416.16 604.88 966.36 769.08
$p$-value << .001
C.2 Annotator Counseling Length of Experience
We examined whether counseling experience is associated with rating behavior. We omitted categories with too few annotators, and chose evaluations by annotators from four experience bands (1–3, 3–5, 5–10, and 10+ years). To ensure comparability, we restricted analysis to the 64 question & response pairs that received at least one evaluation from annotators in each band. For each metric, independence between score category and experience band was tested using the chi-squared test. For Medical Advice and Factual Consistency, “I am not sure” responses were excluded; Medical Advice was treated as a binary per-response indicator.
[…]
Table 9: Mean $\pm$ s.d. scores by annotator counseling experience (1–3, 3–5, 5–10, 10+ years) on the 64 question& response pairs rated by all groups. The bottom row reports $p$-values from P tests of independence between score category and experience band; “I am not sure” responses are excluded for Medical Advice and Factual Consistency; Medical Advice is binary (per response).
Overall Empathy Specificity Medical Advice Factual Consistency Toxicity 1-3 years 3.40 $\pm$ 1.24 3.58 $\pm$ 1.14 3.96 $\pm$ 1.13 0.19 $\pm$ 0.39 3.46 $\pm$ 0.74 1.91 $\pm$ 1.20 3-5 years 3.38 $\pm$ 1.24 3.07 $\pm$ 1.41 3.79 $\pm$ 1.26 0.08 $\pm$ 0.27 3.38 $\pm$ 0.68 1.75 $\pm$ 1.20 5-10 years 3.42 $\pm$ 1.30 3.38 $\pm$ 1.30 3.83 $\pm$ 1.20 0.02 $\pm$ 0.15 3.51 $\pm$ 0.91 1.68 $\pm$ 1.24 10+ years 3.30 $\pm$ 1.22 2.97 $\pm$ 1.35 3.92 $\pm$ 1.17 0.08 $\pm$ 0.27 3.37 $\pm$ 0.87 1.96 $\pm$ 1.34 $p$-value 0.90 0.02 0.83 0.04 0.006 0.27
Appendix D Experiment Details
D.1 Response Generation
We use a zero-shot prompting setup with standardized instructions across models. Decoding parameters are fixed: temperature = 0.7, top-p = 1.0, and a maximum output length of 1024 tokens. To ensure consistency with human responses, we apply rejection sampling to enforce a 250-word limit, regenerating outputs as needed.
We use LABEL:prompt:system-prompt to the LLM to generate a response for each user question, and then provide the original user query directly to generate a response.